# FinCoT: Grounding Chain-of-Thought in Expert Financial Reasoning

### v1

@@ 5.1 Financial Reasoning Performance
We probe transferability by applying each domain–specific FinCoT prompt to the other eight CFA domains and comparing accuracy against Standard Prompting (SP). This reveals how well a blueprint built for one domain generalizes beyond its home context and highlights where expert-aligned structure helps or fails to boost performance outside its intended scope.
We assess cross-domain transfer by applying each domain’s FinCoT blueprint to all CFA domains on Qwen2.5-7B-Instruct and Qwen3-8B-Base. On Qwen2.5-7B-Instruct, the Quantitative Methods, Equity Investments, and Alternative Investments blueprints yield the largest average gains of +3.95 pp, +3.38 pp, and +3.00 pp, respectively [^6] These averages include performance on questions from each blueprint’s own domain.. On Qwen3-8B-Base, the Derivatives, Portfolio Management, and Corporate Issuers blueprints deliver gains of +15.32 pp, +13.66 pp, and +13.54 pp, respectively.
Domains with the smallest positive transfer gains further illustrate this pattern. On Qwen2.5-7B-Instruct, the three lowest nonzero boosts come from Corporate Issuers (+0.90 pp), Economics (+1.60 pp), and Derivatives (+1.72 pp). On Qwen3-8B-Base, the smallest positive transfers are Financial Reporting (+11.31 pp), Economics (+12.15 pp), and Equity Investments (+12.39 pp). In contrast, Alternative Investments transfers minimally (+0.41 pp on Qwen2.5-7B-Instruct; –4.52 pp on Qwen3-8B-Base). We exclude Ethics (<1% of items) from detailed analysis. These trends confirm that broadly quantitative domains generalize well, whereas niche or underrepresented domains show limited or negative transfer.
The resulting cross-domain improvements are plotted in Figure 5, illustrating accuracy gains for each FinCoT prompt in other domains: (a) Qwen2.5-7B-Instruct; (b) Qwen3-8B-Base.
Figure 5: % Impact improvement of each FinCoT across other domains. (a) Qwen2.5 7B; (b) Qwen3-8B-Base.
@@ 5.2 Efficiency Analysis
For an effective LLM deployment, efficiency in terms of scalability and resources must complement accuracy. This section reviews how output verbosity affects accuracy and compares it with Fin-R1, a robust open-source model fine-tuned for financial tasks. Although Fin-R1 is specialized, FinCoT seeks competitive results via prompting, avoiding the need for fine-tuning. These results are visualized in Figure 6, highlighting the comparative performance across prompting strategies.
@@
Figure 6: Comparison of prompt strategies across financial domains. Each point represents a domain-method pair, with position indicating accuracy improvement and domain coverage. Circle size encodes generated tokens; prompt length is listed in Table 5.
Prompt strategies that yield better accuracy with fewer tokens are more desirable in practical applications. FinCoT achieves a higher improvement with moderate token output, striking a balance between the depth of reasoning and the computational cost.
On Qwen3-8B-Base, FinCoT improves accuracy from 78.49% (ST-CoT) to 80.52% (+2.03 pp) while reducing average output length from 3.42k to 0.38 k tokens an $\sim$8.9$\times$ compression. On Qwen3-8B (Thinker), FinCoT matches ST-CoT’s 88.18% accuracy using only 1.23k tokens versus 1.35k ($\sim$1.1$\times$ reduction), indicating that blueprint prompts enable more concise, focused reasoning. Token-coverage analysis reveals three key clusters:
  • FinCoT (All Blueprints): Large green markers in the top-right quadrant reflect overall high domain coverage (80-90%) and notably strong gains (+10-16 pp) with moderate token counts, highlighting wide applicability.
  • Domain-Specific FinCoT: Smaller green circles align diagonally, with core domains achieving 40–60% coverage and +4–8 pp gains. Limited generalizability is reflected in niche domains clustering near the origin.
  • Instruction-Tuned Prompts: Blue/olive markers appear in the lower-coverage region (20-40%), yielding modest gains (+2-6 pp)but with higher token usage, reflecting more verbose internal reasoning.
These patterns highlight FinCoT’s balance between generalization, efficiency, and performance. They suggest that adaptive or hybrid blueprint strategies-leveraging domain relevance-could further improve prompting effectiveness.
@@ 6 Conclusion
We presented FinCoT, a zero-shot prompting framework that embeds carefully crafted domain-specific expert-curated Mermaid within a structured chain-of-thought scaffold. By grounding each reasoning step in domain logic, FinCoT links human financial workflows with embedded expert blueprints without model fine-tuning. The approach is model-agnostic, lightweight to deploy, and produces prompts that are auditable and easy to extend-valuable properties for regulation-heavy domains such as finance.
[…]
@@ Acknowledgments
We extend our gratitude not only to our internal resources but also to the Hatari Team, whose contribution of additional GPU capacity was instrumental in the successful execution of our experiments. Their support significantly augmented our computational capabilities beyond our in-house provisions, facilitating the advancement of our research.
@@ Appendix A Expert Reasoning Blueprints
Step-4: Determine Numerical vs. Conceptual Approach (C) – Numerical (D1): Gather the relevant data (prices, quantities, rates), apply formulae (e.g., $\text{Elasticity}=\frac{\%\Delta Q_{n}}{\%\Delta P}$, $\text{GDP}=C+I+G+(X-M)$ ), and verify assumptions (Investopedia, 2025c). – Conceptual (D2): Construct a narrative explaining cause–effect relationships (e.g., how a monetary-policy change shifts AD or how trade barriers affect capital flows) (Investopedia, 2025b).
@@ B.3 Classify Domain
Appendix C Significance Testing
C.1 Accuracy
Model Baseline Comparison $\Delta$ (pp) 95% CI (pp) $p$-value Significant
Qwen2.5-7B-Instruct SP UST-CoT 6.00 [4.65, 7.56] 0.0000
SP ST-CoT 4.84 [3.59, 6.20] 0.0000
SP FinCoT (All) 4.55 [3.29, 5.91] 0.0000
UST-CoT ST-CoT -1.16 [-1.84, -0.58] 0.0000
UST-CoT FinCoT -1.45 [-2.23, -0.78] 0.0000
ST-CoT FinCoT -0.29 [-0.68, 0.00] 0.1024
Qwen3-8B-Base SP UST-CoT -9.40 [-11.24, -7.66] 0.0000
SP ST-CoT -15.31 [-17.54, -13.18] 0.0000
SP FinCoT (All) -17.35 [-19.77, -15.02] 0.0000
UST-CoT ST-CoT -5.91 [-7.36, -4.55] 0.0000
UST-CoT FinCoT -7.96 [-9.59, -6.30] 0.0000
ST-CoT FinCoT -2.04 [-3.00, -1.26] 0.0000
Qwen3-8B SP UST-CoT 7.95 [6.30, 9.69] 0.0000
SP ST-CoT 6.60 [5.14, 8.14] 0.0000
SP FinCoT (All) 6.70 [5.23, 8.24] 0.0000
UST-CoT ST-CoT -1.35 [-2.13, -0.68] 0.0000
UST-CoT FinCoT -1.26 [-1.94, -0.68] 0.0000
ST-CoT FinCoT 0.10 [0.00, 0.29] 0.7370
Qwen3-8B (Thinker) SP UST-CoT -0.87 [-1.45, -0.39] 0.0002
SP ST-CoT 0.00 [0.00, 0.00] 2.0000
SP FinCoT (All) 0.96 [0.39, 1.55] 0.0002
UST-CoT ST-CoT 0.87 [0.39, 1.45] 0.0002
UST-CoT FinCoT 1.83 [1.07, 2.71] 0.0000
ST-CoT FinCoT 0.96 [0.39, 1.55] 0.0002
Fin-R1 SP UST-CoT -9.49 [-11.34, -7.75] 0.0000
SP ST-CoT -8.62 [-10.37, -6.88] 0.0000
SP FinCoT (All) -10.07 [-11.92, -8.24] 0.0000
UST-CoT ST-CoT 0.87 [0.39, 1.45] 0.0002
UST-CoT FinCoT -0.58 [-1.07, -0.19] 0.0042
ST-CoT FinCoT -1.45 [-2.23, -0.78] 0.0000
Fin-R1 (Strict) SP UST-CoT -1.35 [-2.13, -0.68] 0.0000
SP ST-CoT 0.68 [0.19, 1.26] 0.0016
SP FinCoT (All) -0.10 [-0.29, 0.00] 0.7314
UST-CoT ST-CoT 2.03 [1.26, 2.91] 0.0000
UST-CoT FinCoT 1.26 [0.58, 1.94] 0.0000
ST-CoT FinCoT -0.77 [-1.36, -0.29] 0.0002
@@ C.1 Accuracy
C.2 Average Output Tokens
Model Baseline Comparison $\Delta$ (k) 95% CI (k) $p$-value Significant
Qwen2.5-7B-Instruct SP UST-CoT -0.227 [-0.227, -0.227] 0.0000
SP ST-CoT -0.169 [-0.169, -0.169] 0.0000
SP FinCoT (All) -0.241 [-0.241, -0.241] 0.0000
UST-CoT ST-CoT 0.058 [0.058, 0.058] 0.0000
UST-CoT FinCoT -0.014 [-0.014, -0.014] 0.0000
ST-CoT FinCoT -0.072 [-0.072, -0.072] 0.0000
Qwen3-8B-Base SP UST-CoT -0.584 [-0.584, -0.584] 0.0000
SP ST-CoT 2.522 [2.522, 2.522] 0.0000
SP FinCoT (All) -0.516 [-0.516, -0.516] 0.0000
UST-CoT ST-CoT 3.106 [3.106, 3.106] 0.0000
UST-CoT FinCoT 0.068 [0.068, 0.068] 0.0000
ST-CoT FinCoT -3.038 [-3.038, -3.038] 0.0000
Qwen3-8B SP UST-CoT 0.141 [0.141, 0.141] 0.0000
SP ST-CoT -0.067 [-0.067, -0.067] 0.0000
SP FinCoT (All) 0.048 [0.048, 0.048] 0.0000
UST-CoT ST-CoT -0.208 [-0.208, -0.208] 0.0000
UST-CoT FinCoT -0.093 [-0.093, -0.093] 0.0000
ST-CoT FinCoT 0.115 [0.115, 0.115] 0.0000
Qwen3-8B (Thinker) SP UST-CoT -0.018 [-0.018, -0.018] 0.0000
SP ST-CoT -1.271 [-1.271, -1.271] 0.0000
SP FinCoT (All) -0.168 [-0.168, -0.168] 0.0000
UST-CoT ST-CoT -1.253 [-1.253, -1.253] 0.0000
UST-CoT FinCoT -0.150 [-0.150, -0.150] 0.0000
ST-CoT FinCoT 1.103 [1.103, 1.103] 0.0000
Fin-R1 SP UST-CoT 1.526 [1.526, 1.526] 0.0000
SP ST-CoT 1.338 [1.338, 1.338] 0.0000
SP FinCoT (All) 1.035 [1.035, 1.035] 0.0000
UST-CoT ST-CoT -0.188 [-0.188, -0.188] 0.0000
UST-CoT FinCoT -0.491 [-0.491, -0.491] 0.0000
ST-CoT FinCoT -0.303 [-0.303, -0.303] 0.0000
Fin-R1 (Strict) SP UST-CoT -0.099 [-0.099, -0.099] 0.0000
SP ST-CoT 0.023 [0.023, 0.023] 0.0000
SP FinCoT (All) -0.116 [-0.116, -0.116] 0.0000
UST-CoT ST-CoT 0.121 [0.121, 0.121] 0.0000
UST-CoT FinCoT -0.017 [-0.017, -0.017] 0.0000
ST-CoT FinCoT -0.138 [-0.138, -0.138] 0.0000

### v2

@@ 5.1 Financial Reasoning Performance
This section examines how pretrained and finance-specific models respond to structured prompting with FinCoT. Each domain-specific blueprint is applied across all CFA domains to evaluate its transferability, and accuracy differences relative to SP are measured. Figure 4(a) and 4(b) visualize results for Qwen3-8B-Base (pretrained) and Fin-o1-8B (finance-specific), while Table 1 provides a comprehensive summary of overall model-level accuracy. On Qwen3-8B-Base, FinCoT consistently improves performance. Prompts from quantitative domains such as Derivatives, Portfolio Management, and Corporate Issuers yield average gains exceeding +13 pp. The blueprint structure provides inductive guidance that enhances decomposition, formula selection, and financial term alignment. In several domains, Qwen3-8B-Base with FinCoT matches or surpasses the SP baseline of Fin-o1-8B, despite the absence of any task-specific training.
On Fin-o1-8B, gains from FinCoT are more modest, typically within the +1–4 pp range. Minor declines appear in some domains (e.g., Fixed Income, Equity Investments), suggesting that additional scaffolding may interfere with optimized internal reasoning acquired during fine-tuning. Structured prompts may over-specify solutions or reduce instruction-following flexibility.
These findings highlight FinCoT’s complementary role. For pretrained models, FinCoT acts as a lightweight yet effective augmentation layer at inference time, reducing the performance gap with fine-tuned models. For already fine-tuned models, careful prompt selection or adaptation may be necessary to preserve existing reasoning strengths without introducing conflict. A broader breakdown of FinCoT performance across additional models is provided in Appendix F, where radar plots illustrate domain-wise patterns and complement the main analysis.
Figure 5: Comparison of prompt strategies across financial domains. Each point represents a domain-method pair, with position indicating accuracy improvement and domain coverage. Circle size encodes generated tokens.
@@ 5.2 Efficiency Analysis
Effective deployment of foundation models in financial settings requires not only accuracy but also efficiency in output length and computational cost. This section investigates the verbosity and accuracy trade-offs across prompting strategies. While Fin-o1-8B, DianJin-R1-7B, and Fin-R1 (7B) are fine-tuned models optimized for financial reasoning, FinCoT offers a prompting-based alternative that achieves competitive performance without additional training. Comparative accuracy is illustrated in Figure 5, and a detailed breakdown of output lengths across models and domains is provided in Appendix D (Table 2) to support efficiency analysis.
@@
In practical deployments, prompting strategies that achieve high accuracy with minimal output length are especially desirable. This section compares the token efficiency of FinCoT with standard prompting strategies, focusing on general-purpose and financial-specific models.
For general-purpose models such as Qwen3-8B-Base and Qwen3-8B (Thinker), FinCoT substantially reduces output length while preserving or improving accuracy. On Qwen3-8B-Base, FinCoT improves accuracy from 78.49% (ST-CoT) to 80.52% (+2.03 pp), while reducing average output from 3.42k to 0.38k tokens, an approximately 8.9$\times$ compression. Similarly, on Qwen3-8B (Thinker), FinCoT maintains parity with ST-CoT at 88.18% accuracy, using 1.23k tokens compared to 1.35k, a $\sim$1.1$\times$ reduction. These results support the view that FinCoT’s structured blueprints enable more concise, focused reasoning in general-purpose models. Among financial-specific models such as DianJin-R1-7B and Fin-o1-8B, FinCoT shows minimal or negative improvement and little benefit in output compression. This aligns with the idea that fine-tuned models already internalize domain-specific reasoning. In particular, Fin-o1-8B illustrates that excessive prompt scaffolding may interfere with latent reasoning, reducing effectiveness and leading to what can be described as overthinking that undermines transfer.
Three behavioral zones emerge from these results. First, models like Gemma-3-12B-IT and Qwen3-8B-Base show high mean improvement and strong cross-domain transfer, reflecting effective generalization. Second, models such as Qwen2.5-7B-Instruct and Qwen3-8B display noticeable cross-domain transfer with lower mean improvement, suggesting limited benefit. Third, models like Fin-o1-8B and Qwen3-8B (Thinker) exhibit low cross-domain adaptability and minimal performance lift, with signs that overly detailed prompting may interfere with internal reasoning.
These findings highlight the importance of aligning the prompting strategy with a model’s pretraining or fine-tuning characteristics to optimize both performance and efficiency. With limited supervision, FinCoT offers a transparent, cost-effective alternative for enhancing financial reasoning.
@@ 6 Conclusion
We presented FinCoT, a zero-shot prompting framework that embeds expert-curated Mermaid diagrams within structured chain-of-thought scaffolds. By grounding reasoning in domain logic, FinCoT bridges human financial workflows with LLM outputs, without model fine-tuning. While broadly applicable, FinCoT yields strong gains for general models (e.g., Qwen3-8B-Base), but more modest or negative effects for instruction- or finance-specific models (e.g., Qwen-Thinker, Fin-o1, DianJin-R1), where added structure may interfere with learned reasoning.
Relative to SP, FinCoT improves Qwen3-8B-Base by +17.33 pp and Fin-R1 by +10.08 pp ($p<0.001$), outperforming some fine-tuned models. In contrast, instruction-tuned models like Qwen3-8B (Thinker) sometimes favor UST-CoT. Cross-domain results show blueprints from quantitative fields transfer best (+27.3 pp on Qwen3-8B-Base). FinCoT also reduces token output by up to 8×8$\times$8 ×, offering interpretable, efficient prompting for regulated financial applications. Ultimately, FinCoT suggests that with meticulous prompt design, even general-purpose LLMs can approach the reasoning quality of fine-tuned financial experts in complex decision-making tasks.
@@ Acknowledgments
We extend our gratitude to the Hatari Team for generously providing additional GPU resources, which significantly enhanced our computational capacity and enabled the successful execution of large-scale experiments. We also thank Tuntai Boriboonthana (CFA Level-III holder) from InnovestX for his expert contributions in verifying blueprint fidelity and conducting random audits to ensure domain classification consistency in GPT-4o outputs. Their support and insights were invaluable to the development and evaluation of this work.
@@ Appendix A Expert Reasoning Blueprints
Step-4: Determine Numerical vs. Conceptual Approach (C) – Numerical (D1): Gather the relevant data (prices, quantities, rates), apply formulae (e.g., $\text{Elasticity}=\frac{\%\Delta Q_{n}}{\%\Delta P}$, $)\text{GDP}=C+I+G+(X-M)$, and verify assumptions (Investopedia, 2025c). – Conceptual (D2): Construct a narrative explaining cause–effect relationships (e.g., how a monetary-policy change shifts AD or how trade barriers affect capital flows) (Investopedia, 2025b).
@@ B.3 Classify Domain
Appendix C Domain Distribution
Appendix D Average Output Tokens
Prompt Average Output Tokens (k)
Qwen2.5-7B
Qwen2.5-7B
Instruct
Qwen3-8B
Base
Qwen3-8B
Gemma-3-12B
IT
Qwen3-8B
Fin-R1
DianJin-R1
7B
Fin-o1-8B
SP 0.45 0.05* 0.89 0.32 0.27* 1.52 0.88 2.18 0.46*
UST-CoT 0.48 0.28 0.31* 0.46 0.39 1.50 0.58* 2.28 0.53
ST-CoT 0.39* 0.22 3.42 0.25* 0.31 1.35 2.22 7.20 0.58
FinCoT (All Blueprints) 2.22 0.29 0.38 0.36 0.32 1.23* 1.92 1.60* 0.79
Domain-wise performance of FinCoT
FinCoT (Economics) 0.36 0.38 0.99 0.39 0.38 1.25 2.01 12.65 0.76
FinCoT (FixedIncome) 0.42 0.27 4.55 0.30 0.32 1.24 2.31 8.31 0.81
FinCoT (Quant.Meth.) 0.48 0.27 3.07 0.31 0.35 1.22 2.17 8.60 0.80
FinCoT (EquityInvest.) 0.32 0.31 7.18 0.37 0.34 1.19 2.16 10.07 0.78
FinCoT (Port.Manage.) 0.38 0.26 0.56 0.30 0.33 1.20 2.14 9.46 0.79
FinCoT (Derivatives) 0.36 0.30 0.42 0.39 0.34 1.24 2.05 5.54 0.81
FinCoT (Fin. Reporting) 0.46 0.28 0.93 0.33 0.34 1.19 2.13 8.76 0.73
FinCoT (Alter.Invest.) 0.47 0.26 0.50 0.38 0.34 1.23 2.16 11.53 0.77
FinCoT (Corp.Issuers) 0.52 0.26 1.18 0.32 0.33 1.16 2.08 11.37 0.82
Table 2: Comparison of prompting techniques: average output token length (k) across models. Bold values highlight the prompt variant that uses the least tokens for each model. (*) Indicates that the change in average output token count among the model-level prompt variants is statistically significant ($p<0.05$) based on paired bootstrap testing; domain-specific rows are not tested for significance.
Appendix E Significance Testing
E.1 Accuracy
Model Baseline Comparison $\Delta$ (pp) 95% CI (pp) $p$-value Significant Qwen2.5-7B SP UST-CoT -13.76 [-15.89, -11.63] 0.0000 SP ST-CoT -16.27 [-18.60, -14.05] 0.0000 SP FinCoT (All) -7.94 [ -9.59, -6.30] 0.0000 UST-CoT ST-CoT -2.52 [ -3.49, -1.65] 0.0000 UST-CoT FinCoT (All) 5.81 [ 4.46, 7.27] 0.0000 ST-CoT FinCoT (All) 8.33 [ 6.69, 9.98] 0.0000 Qwen2.5-7B-Instruct SP UST-CoT 6.00 [ 4.65, 7.56] 0.0000 SP ST-CoT 4.84 [ 3.59, 6.20] 0.0000 SP FinCoT (All) 4.55 [ 3.29, 5.91] 0.0000 UST-CoT ST-CoT -1.16 [-1.84, -0.58] 0.0000 UST-CoT FinCoT (All) -1.45 [-2.23, -0.78] 0.0000 ST-CoT FinCoT (All) -0.29 [-0.68, 0.00] 0.1024 Qwen3-8B-Base SP UST-CoT -9.40 [-11.24, -7.66] 0.0000 SP ST-CoT -15.31 [-17.54,-13.18] 0.0000 SP FinCoT (All) -17.35 [-19.77,-15.02] 0.0000 UST-CoT ST-CoT -5.91 [-7.36, -4.55] 0.0000 UST-CoT FinCoT (All) -7.96 [-9.59, -6.30] 0.0000 ST-CoT FinCoT (All) -2.04 [-3.00, -1.26] 0.0000 Qwen3-8B SP UST-CoT 7.95 [ 6.30, 9.69] 0.0000 SP ST-CoT 6.60 [ 5.14, 8.14] 0.0000 SP FinCoT (All) 6.70 [ 5.23, 8.24] 0.0000 UST-CoT ST-CoT -1.35 [-2.13, -0.68] 0.0000 UST-CoT FinCoT (All) -1.26 [-1.94, -0.68] 0.0000 ST-CoT FinCoT (All) 0.10 [ 0.00, 0.29] 0.7370 Qwen3-8B (Thinker) SP UST-CoT -0.87 [-1.45, -0.39] 0.0002 SP ST-CoT 0.00 [ 0.00, 0.00] 2.0000 SP FinCoT (All) 0.96 [ 0.39, 1.55] 0.0002 UST-CoT ST-CoT 0.87 [ 0.39, 1.45] 0.0002 UST-CoT FinCoT (All) 1.83 [ 1.07, 2.71] 0.0000 ST-CoT FinCoT (All) 0.96 [ 0.39, 1.55] 0.0002 Gemma-3-12B-IT SP UST-CoT -24.999 [-27.71, -22.38] 0.0000 SP ST-CoT -23.934 [-26.55, -21.32] 0.0000 SP FinCoT (All) -22.765 [-25.39, -20.16] 0.0000 UST-CoT ST-CoT 1.065 [ 0.48, 1.74] 0.0002 UST-CoT FinCoT (All) 2.234 [ 1.36, 3.20] 0.0000 ST-CoT FinCoT (All) 1.169 [ 0.58, 1.84] 0.0000 Fin-R1 SP UST-CoT -9.49 [-11.34, -7.75] 0.0000 SP ST-CoT -8.62 [-10.37, -6.88] 0.0000 SP FinCoT (All) -10.07 [-11.92,-8.24] 0.0000 UST-CoT ST-CoT 0.87 [ 0.39, 1.45] 0.0002 UST-CoT FinCoT (All) -0.58 [-1.07, -0.19] 0.0042 ST-CoT FinCoT (All) -1.45 [-2.23, -0.78] 0.0000 Dianjin-R1-7B SP UST-CoT 10.662 [ 8.82, 12.60] 0.0000 SP ST-CoT 9.594 [ 7.75, 11.43] 0.0000 SP FinCoT (All) -1.361 [ -2.13, -0.68] 0.0000 UST-CoT ST-CoT -1.068 [ -1.74, -0.48] 0.0000 UST-CoT FinCoT (All) -12.023 [-14.05, -10.08] 0.0000 ST-CoT FinCoT (All) -10.956 [-12.89, -9.01] 0.0000 Fino1-8B SP UST-CoT 0.294 [ 0.00, 0.68] 0.0984 SP ST-CoT 1.264 [ 0.58, 1.94] 0.0000 SP FinCoT (All) 2.429 [ 1.55, 3.39] 0.0000 UST-CoT ST-CoT 0.970 [ 0.39, 1.55] 0.0000 UST-CoT FinCoT (All) 2.134 [ 1.26, 3.00] 0.0000 ST-CoT FinCoT (All) 1.165 [ 0.58, 1.84] 0.0000
@@ E.1 Accuracy
E.2 Average Output Tokens
Model Baseline Comparison $\Delta$ (k) 95% CI (k) $p$-value Significant Qwen2.5-7B SP UST-CoT -0.09867 [-0.09867, -0.09867] 0.0000 SP ST-CoT 0.02273 [ 0.02273, 0.02273] 0.0000 SP FinCoT (All) -0.11555 [-0.11555, -0.11555] 0.0000 UST-CoT ST-CoT 0.12140 [ 0.12140, 0.12140] 0.0000 UST-CoT FinCoT (All) -0.01688 [-0.01688, -0.01688] 0.0000 ST-CoT FinCoT (All) -0.13828 [-0.13828, -0.13828] 0.0000 Qwen2.5-7B-Instruct SP UST-CoT -0.227 [-0.227, -0.227] 0.0000 SP ST-CoT -0.169 [-0.169, -0.169] 0.0000 SP FinCoT (All) -0.241 [-0.241, -0.241] 0.0000 UST-CoT ST-CoT 0.058 [ 0.058, 0.058] 0.0000 UST-CoT FinCoT (All) -0.014 [-0.014, -0.014] 0.0000 ST-CoT FinCoT (All) -0.072 [-0.072, -0.072] 0.0000 Qwen3-8B-Base SP UST-CoT -0.584 [-0.584, -0.584] 0.0000 SP ST-CoT 2.522 [ 2.522, 2.522] 0.0000 SP FinCoT (All) -0.516 [-0.516, -0.516] 0.0000 UST-CoT ST-CoT 3.106 [ 3.106, 3.106] 0.0000 UST-CoT FinCoT (All) 0.068 [ 0.068, 0.068] 0.0000 ST-CoT FinCoT (All) -3.038 [-3.038, -3.038] 0.0000 Qwen3-8B SP UST-CoT 0.141 [ 0.141, 0.141] 0.0000 SP ST-CoT -0.067 [-0.067, -0.067] 0.0000 SP FinCoT (All) 0.048 [ 0.048, 0.048] 0.0000 UST-CoT ST-CoT -0.208 [-0.208, -0.208] 0.0000 UST-CoT FinCoT (All) -0.093 [-0.093, -0.093] 0.0000 ST-CoT FinCoT (All) 0.115 [ 0.115, 0.115] 0.0000 Qwen3-8B (Thinker) SP UST-CoT -0.018 [-0.018, -0.018] 0.0000 SP ST-CoT -1.271 [-1.271, -1.271] 0.0000 SP FinCoT (All) -0.168 [-0.168, -0.168] 0.0000 UST-CoT ST-CoT -1.253 [-1.253, -1.253] 0.0000 UST-CoT FinCoT (All) -0.150 [-0.150, -0.150] 0.0000 ST-CoT FinCoT (All) 1.103 [ 1.103, 1.103] 0.0000 Gemma-3-12B-IT SP UST-CoT 0.11985 [ 0.11985, 0.11985] 0.0000 SP ST-CoT 0.03661 [ 0.03661, 0.03661] 0.0000 SP FinCoT (All) 0.04523 [ 0.04523, 0.04523] 0.0000 UST-CoT ST-CoT -0.08324 [-0.08324, -0.08324] 0.0000 UST-CoT FinCoT (All) -0.07462 [-0.07462, -0.07462] 0.0000 ST-CoT FinCoT (All) 0.00862 [ 0.00862, 0.00862] 0.0000 Fin-R1 SP UST-CoT 1.526 [ 1.526, 1.526] 0.0000 SP ST-CoT 1.338 [ 1.338, 1.338] 0.0000 SP FinCoT (All) 1.035 [ 1.035, 1.035] 0.0000 UST-CoT ST-CoT -0.188 [-0.188, -0.188] 0.0000 UST-CoT FinCoT (All) -0.491 [-0.491, -0.491] 0.0000 ST-CoT FinCoT (All) -0.303 [-0.303, -0.303] 0.0000 Dianjin-R1-7B SP UST-CoT 0.10023 [ 0.10023, 0.10023] 0.0000 SP ST-CoT 5.02159 [ 5.02159, 5.02159] 0.0000 SP FinCoT (All) -0.57669 [-0.57669, -0.57669] 0.0000 UST-CoT ST-CoT 4.92136 [ 4.92136, 4.92136] 0.0000 UST-CoT FinCoT (All) -0.67692 [-0.67692, -0.67692] 0.0000 ST-CoT FinCoT (All) -5.59828 [-5.59828, -5.59828] 0.0000 Fino1-8B SP UST-CoT 0.06788 [ 0.06788, 0.06788] 0.0000 SP ST-CoT 0.11765 [ 0.11765, 0.11765] 0.0000 SP FinCoT (All) 0.33273 [ 0.33273, 0.33273] 0.0000 UST-CoT ST-CoT 0.04977 [ 0.04977, 0.04977] 0.0000 UST-CoT FinCoT (All) 0.26485 [ 0.26485, 0.26485] 0.0000 ST-CoT FinCoT (All) 0.21508 [ 0.21508, 0.21508] 0.0000